An integrating factor of (x / 2) dx + ( x^2y + 8y) dy = 0 is a. My b. eyn2) c^{x^2} None of the above 42y^2)
Answer #1


X dx+(x+y+8y)dy = 0 This is in the form Mdx + Ndy = 0 Here, M = and N = x+y+8y 2 ам ƏN 0, = 29+0) = 2x OM since, ON It is not